# EDUMATH: Generating Standards-aligned Educational Math Word Problems

### v1

@@ 3.2 Synthetic Data Generation and Expert Annotation
While our labeled dataset provides MWPs for prompting, it does not have enough MWPs for training a standards-aligned educational MWP generator given its small size. As shown in Appendix C, most of the 23 VA SOL combinations in the labeled ASDIV subset also only have a few MWPs they are associated with. Further review of VA SOLs also revealed another 15 combinations of standards that can be addressed by MWPs but are not in the labeled ASDIV subset. To address these gaps, we generated a large synthetic dataset (n = 3,012) and evaluated it with teachers. We generated this data with Llama 3.3 70B IT Meta using the prompt in Appendix I.4. We generated roughly 90 samples per VA SOL combination that is not in our labeled ASDIV subset and roughly 80 samples per VA SOL combination that is in the labeled subset but has less than 100 samples to ensure sufficient coverage of all relevant VA SOLs while emphasizing those that are not already included.
Next, teachers annotated our generated MWPs through Prolific. Annotators (n = 1,372) identified as US teachers using Prolific screening criteria and were paid \$12/hour. Annotators labeled MWPs for solvability, accuracy, educational appropriateness, and standards alignment using the directions in Appendix A. See Appendix A for annotator demographics and time spent per MWP. Every MWP was annotated twice; if the annotators disagreed on any of the criteria, the MWP was annotated a third time. We assigned final labels based on majority vote. We labeled a question as meeting all criteria (MaC) if it was labeled as solvable, accurate, educationally appropriate, and standards-aligned.
The first two annotators agreed on solvability $90.1\pm 1.2\%$ of the time, accuracy $76.6\pm 1.5\%$ of the time, educational appropriateness $77.5\pm 1.5\%$ of the time, standards alignment $75.3\pm 1.6\%$ of time, and MaC $65.5\pm 1.7\%$ of the time. The agreement rate for solvability and appropriateness is higher than that reported in Christ et al. (2024), while the agreement rate for accuracy and MaC is lower. The lower accuracy agreement rate could be because accuracy in this work evaluates both final answer and intermediate reasoning accuracy as well as solution readability, whereas Christ et al. (2024) only evaluated final answer accuracy. The lower MaC agreement rate could be because it is based on agreement across four metrics instead of three, as was the case in Christ et al. (2024). However, to ensure only the highest quality data are included in our synthetic training dataset, we had each MWP annotated by Gemma 3 27B IT using few-shot examples and the same directions given to annotators (see Appendix I.5 for the prompt). We flipped the label for any MWP humans identified as MaC but Gemma identified as not MaC, but retained the label for any MWP humans identified as not MaC but Gemma identified as MaC.
@@ 3.3 Standards-Targeted Educational Math Dataset
Our annotation process identified 1,552 MWPs that MaC, and we compile these MWPs along with the labeled ASDIV subset into a new dataset we call Standards-Targeted Educational Math (STEM). Table 1 shows STEM’s advantages over other grade school MWP datasets. Notably, STEM is the only dataset annotated by teachers for math standards and with solutions designed to be readable for students and teachers, making it best suited for training a standards-aligned educational MWP generator. On other metrics, STEM is similar to existing datasets. STEM’s question length is in the middle of the range for grade school math datasets, while its average readability is slightly higher than these datasets, suggesting it may be a little more difficult to read. STEM’s BERTScore is similar to that of human-written datasets, suggesting its MWPs are similar. Overall, these metrics highlight that STEM is similar to human-written data but tailored to standards-aligned educational MWP generation.
@@ 3.4 Finetuning on Annotated Data
To explore training a standards-aligned educational MWP generator and validate the quality of our labeled data, we use it to finetune a mid-size open LLM, Gemma 3 12B IT Team et al. (2025). We first conduct Supervised Finetuning (SFT) using STEM. Next, we conduct Kahneman-Tversky Optimization (KTO) Ethayarajh et al. (2024) on our SFT model with our annotated dataset given MaC labels are binary. Next, we train a ModernBERT text classifier Warner et al. (2024) using our annotation data to label outputs for MaC that reaches 79% accuracy and a 0.861 AUC-ROC on a held-out test set. We then use this classifier on top of our trained KTO model to filter outputs to only those labeled as high quality, resulting in the model we call EDUMATH 12B. In Section 4, we also stack our classifier on top of Qwen 3 30B Yang et al. (2025) to create EDUMATH 30B. Appendix B.3 shows ablations for EDUMATH 12B’s training using our Gemma 3 27B IT annotator. Both finetuning stages and automatic filtering using our text classifier significantly improve model performance, showing the quality of our data. See Appendix B for further training details and hyperparameters.
@@ 4 Evaluating EDUMATH Models
Next, we rigorously evaluate both EDUMATH models’ generated MWPs using our evaluation criteria and automatic evaluations. We compare 1,000 MWPs from each EDUMATH model to those of open and closed models. For closed models, we evaluate 250 MWPs from each of GPT-4o OpenAI et al. (2024), GPT-4.1 OpenAI (2025), and GPT-4.5 OpenAI (2025). We use a smaller set of MWPs for closed models due to API cost. For open models, we evaluate 1,000 MWPs from Gemma 3 12B IT Team et al. (2025), Gemma 3 27B IT Team et al. (2024), Qwen 3 30B IT Yang et al. (2025), and Qwen 3 235B IT Yang et al. (2025), which all represent highly capable open baselines of various sizes. We prompt each model using 8-shot examples from STEM for a given math standard and ask it to create a new MWP based on the same math standard (see Appendix I.4 for prompting details).
@@
Table 2: Comparing LLMs for standards-aligned educational MWP generation. All metrics are averages over 1,000 MWPs/model except for GPT models (they include 250 MWPs due to cost). PPL is perplexity, BF1 is BERTScore F1, ASDIV BF1 compares each model’s questions to the ASDIV 3rd-5th grade subset, length is average token length, readability is a composite of 8 metrics, and MaC is meets all criteria. Bold indicates the lowest PPL and best MaC performance and a * indicates the difference between the best open MaC performance and second best open MaC performance is significant at the p < 0.01 level. Standard errors, where appropriate, are in parentheses.
@@ 4.1 Main Results
Using our Gemma 3 27B IT annotator, we evaluated each model’s MWPs for solvability, accuracy, educational appropriateness, and standards alignment. To reduce LLM calls, we annotate for all criteria simultaneously and have the LLM output if the MWP MaC along with a CoT explaining its reasoning (see Appendix I.5 for an example). To validate our labels, our research team member with K-12 teaching experience and undergraduate education student labeled 100 randomly selected samples from our evaluation set. The annotators agreed with each other 76% of the time on MaC and had an average agreement rate of 75% with Gemma 3 27B IT for MaC. Given the agreement rates are nearly identical and higher than those reported among teachers in Section 3.2, we are confident in our labels.
Table 2 shows the MaC rate for each model. Our 30B model outperforms all open and closed baselines in MaC, while our 12B model outperforms Gemma 3 27B IT and nearly matches the performance of Qwen 3 30B IT, two much larger and more capable models. The table also shows performance gaps between open and closed models and between smaller open and larger open models that our data/models eliminate or nearly eliminate. These results suggest our data are sufficient for enabling open models to meet SOTA performance.
@@
Table 2 also shows PPL, BERTScore, and the average question and solution length and readability for each model’s MWPs. Our 12B model has the lowest PPL among all models and our 30B model has the lowest PPL among larger models, suggesting their MWPs are high quality. Our 12B model’s BERTScore is most similar to ASDIV’s within-dataset BERTScore among all models, while our 30B model ties with Qwen 3 235B IT in having a BERTScore closest to ASDIV’s among the open models excluding EDUMATH 12B. The BERTScores of our models’ MWPs compared to ASDIV is roughly equivalent to ASDIV’s within-dataset BERTScore, suggesting their MWPs are similar in diversity to human-written MWPs, while the MWPs from other models have higher BERTScores for this comparison. Our models’ average question lengths are most similar to those of STEM and ASDIV, suggesting their complexity, as measured by this proxy, is similar to teacher-written MWPs. Our models have the shortest solutions among open models, though the closed baselines have the shortest solutions among all models. Shorter solutions are preferable to longer, more verbose ones that may be daunting to read.
For readability, all models output questions with readability similar to STEM except that Qwen 3 235B IT, GPT 4.1, and EDUMATH 12B slightly exceed STEM’s readability. For solution readability, all open models’ solutions have lower readability than STEM, while closed models exceed STEM’s readability, denoting a strength of open models. Overall, Table 2 suggests the EDUMATH models’ MWPs are high quality and similar to human-written MWPs with the exception that EDUMATH 12B’s MWPs are slightly harder to read.
@@
Although there are 38 math standard combinations, they can be grouped into combinations of one or more of 8 math topics: addition/subtraction, multiplication/division, fractions, decimals, perimeter/area/volume, patterns, elapsed time, and measurement conversion. Exploring MaC rates by topic can highlight the topics each model struggles with and succeeds at. Table 3 shows this comparison, while Appendix D shows further analysis of common error types for each model. EDUMATH 12B performs best at time problems and struggles the most with fraction problems, while EDUMATH 30B performs best at conversion problems and struggles the most with multiplication/division problems. EDUMATH 12B and the Gemma models have larger gaps in performance between the topic they perform best and worse at than the other models. EDUMATH 30B is the only model for which performance on each topic is above 90%, further highlighting its SOTA performance.
@@ 5 Evaluating MWPs with Students
While we show our models’ MWPs are high-quality and can be customized to student interests (see Appendix E), it is critical to ensure students perform similarly on their MWPs relative to human-written MWPs and to identify if students prefer their customized MWPs to non-customized MWPs. To test this, 3rd-5th graders solved a worksheet with one MWP generated by EDUMATH 30B and aligned with topics they are interested in and one written by humans, either from existing data (K5Learning or the ASDIV 3rd-5th grade subset) or written by our research team member with K-12 teaching experience if a worksheet’s math topic was not captured in existing data (see Appendix F.3 for an example). We also asked students which MWP they preferred and why. We tested these worksheets in two experimental settings at two elementary schools (see Appendix F.2 for school demographics). All worksheets were screened by students’ teachers to ensure the MWPs were appropriate and reasonable. At the first school, 4th graders (n = 82) solved one worksheet for each of four weeks where both the LLM-generated and human-written MWP were the same for all students. Each MWP was aligned with a standard students were currently learning and the LLM-generated MWP was based on a topic teachers identified as interesting to students. At the second school, 3rd-5th graders (n = 12) receiving services through a math interventionist completed one or two worksheets with the same human-written MWP for every student in each grade level and a customized LLM-generated MWP for each student in that grade level, depending on how many standards they were learning (see Appendix F.2 for details). To inform the customized MWPs, the math interventionist surveyed students to identify topics they were interested in, resulting in a high level of customization for each student. These two experimental conditions allow us to test differential effects of customization at the class and individual level, while also testing our MWPs with students across the grade levels we study.
Through this experiment (see Appendix F.1 for full results), we found students performed similarly on human-written and LLM-generated MWPs, further suggesting EDUMATH 30B’s MWPs are human quality. Students preferred the LLM-generated MWPs over the human-written MWPs in both conditions. In School 2 where customization was at the individual level, every student but one preferred the LLM-generated MWPs. Students most often reported liking the LLM-generated MWP because they liked the topic it was about. These results offer compelling evidence that LLMs can support elementary math education by providing customized MWPs students find exciting.
@@ 6 Conclusion
We explore standards-aligned educational MWP generation and create the first dataset to train models for this task and the only one verified by teachers to have readable solutions. We also conduct an evaluation of the performance of existing models at this task and fully release our annotations to aid the research community. We demonstrate the quality of STEM and our annotated data by using them to develop the EDUMATH models, the first standards-aligned educational MWP generators. Our evaluations show EDUMATH 12B matches the performance of larger and more capable open models, while EDUMATH 30B sets a new SOTA. We also conduct the first study of custom LLM-generated MWPs with grade school students, finding they perform similarly on our MWPs to those that are human-written, while preferring our MWPs. Our findings suggest that LLMs are poised to greatly contribute to K-12 education by directly generating customized practice problems for students to reduce teacher burdens. Future research should develop models and data to generate problems for other grade levels and types of math problems.
@@
Appendix A Annotator Directions and Demographics
Figures 2, 3, 4, and 5 display the directions presented to annotators for annotating MWPs for solvability, accuracy, educational appropriateness, and standards alignment, respectively. If an annotator answered "No" to solvability or any other criteria, they were instructed to select "NA" for the remaining criteria given that a failure to meet one of the criteria automatically means a MWP does not MaC. This saved the annotators time, as they did not need to unnecessarily annotate the remaining criteria. Table 4 displays annotator demographics and average time spent per MWP. Our human annotation study was approved by the institutional review board at our host university.

### v2

@@ 3.2 Synthetic Data Generation and Expert Annotation
While our labeled dataset provides MWPs for prompting, it is not sufficient for training a standards-aligned educational MWP generator given its small size. As shown in Appendix D, most of the 23 VA SOL combinations in the labeled dataset only have a few MWPs they are associated with. Further review of VA SOLs revealed another 15 combinations of standards that can be addressed by MWPs but are not in the labeled dataset. To address these gaps, we generated a large synthetic dataset (n = 3,012) and evaluated it with teachers. We generated these data with Llama 3.3 70B IT Meta using the prompt in Appendix J.4. We generated roughly 90 samples per VA SOL combination that is not in our labeled dataset and roughly 80 samples per VA SOL combination that is in the labeled dataset but has less than 100 samples to ensure coverage of all relevant VA SOLs while emphasizing those that are not already included.
Next, teachers annotated generated MWPs through Prolific. Annotators (n = 1,372) identified as US teachers and were paid \$12/hour. Annotators labeled MWPs for solvability, accuracy, educational appropriateness, and standards alignment using the directions in Appendix B. See Appendix B for demographics and time spent per MWP. Each MWP was annotated twice; if the annotators disagreed on any criteria, the MWP was annotated a third time. We assigned final labels based on majority vote. We labeled a question as meeting all criteria (MaC) if it was labeled as solvable, accurate, educationally appropriate, and standards-aligned.
The first two annotators agreed on solvability $90.1\pm 1.2\%$ of the time, accuracy $76.6\pm 1.5\%$ of the time, educational appropriateness $77.5\pm 1.5\%$ of the time, standards alignment $75.3\pm 1.6\%$ of time, and MaC $65.5\pm 1.7\%$ of the time. The agreement rate for solvability and appropriateness is higher than that reported in Christ et al. (2024), while the agreement rate for accuracy and MaC is lower. The lower accuracy agreement rate could be because accuracy in this work evaluates both final answer and intermediate reasoning accuracy as well as solution readability, whereas Christ et al. (2024) only evaluated final answer accuracy. The lower MaC agreement rate could be because it is based on agreement across four metrics instead of three like Christ et al. (2024). However, to ensure only the highest quality data are included in our training dataset, we had each MWP annotated by Gemma 3 27B IT using few-shot examples and the same directions given to annotators (see Appendix J.5 for the prompt). We flipped the label for MWPs teachers marked as MaC but Gemma 3 27B IT did not, ensuring our final dataset contains only the highest-quality MWPs and reducing potential human labeling errors. We retained the label for any MWP humans identified as not MaC and Gemma 3 27B IT identified as MaC to rely on teachers’ expertise in cases where MWPs do not MaC due to a nuance only human experts can identify.
@@ 3.3 Standards-Targeted Educational Math Dataset
Our annotation process identified 1,552 MWPs that MaC, and we compile these MWPs along with the labeled ASDIV subset into a new dataset we call Standards-Targeted Educational Math (STEM). Table 1 shows STEM’s advantages over other grade school MWP datasets. Notably, STEM is the only dataset annotated by teachers for math standards and with solutions designed to be readable for students and teachers, making it best suited for training a standards-aligned educational MWP generator. On other metrics, STEM is similar to existing datasets. STEM’s question length is in the middle of the range for grade school math datasets, while its average readability is slightly higher than these datasets. STEM’s BERTScore is similar to that of human-written datasets, suggesting its MWPs are similar. Overall, these metrics highlight that STEM is similar to human-written data but tailored to standards-aligned educational MWP generation.
@@ 3.4 Finetuning on Annotated Data
To validate our dataset’s quality, we use it to finetune a mid-size open LLM, Gemma 3 12B IT Team et al. (2025). We first conduct Supervised Finetuning (SFT) using STEM. Next, we conduct Kahneman-Tversky Optimization (KTO) Ethayarajh et al. (2024) on our SFT model with our annotated data given MaC labels are binary. Next, we train a ModernBERT text classifier Warner et al. (2024) using our annotation data to label MWPs for MaC that reaches 79% accuracy and a 0.861 AUC-ROC on a held-out test set. We then use this classifier on top of our trained KTO model to filter MWPs to only those labeled as high quality, resulting in EDUMATH 12B. [^2] Using a post-processing classifier is akin to how proprietary models reach SOTA performance by using classifiers or multiple models (e.g., OpenAI (2025a)) We also stack our classifier on top of Qwen 3 30B Yang et al. (2025) to create EDUMATH 30B. Appendix C.3 shows ablations for EDUMATH 12B’s training. Both finetuning stages and automatic filtering using our text classifier significantly improve model performance, showing the quality of our data. See Appendix C for further training details and hyperparameters.
@@ 4 Evaluating EDUMATH Models
Next, we rigorously evaluate both EDUMATH models using our evaluation criteria and automatic evaluations. We compare 1,000 MWPs from each EDUMATH model to those of open and closed models to see if small models trained on our high-quality data outperform larger, general purpose models. For closed models, we evaluate 250 MWPs from each of GPT-4o OpenAI et al. (2024), GPT-4.1 OpenAI (2025b), and GPT-4.5 OpenAI (2025b). We use a smaller set of MWPs for closed models due to API cost. For open models, we evaluate 1,000 MWPs from Gemma 3 12B IT Team et al. (2025), Gemma 3 27B IT Team et al. (2024), Qwen 3 30B IT Yang et al. (2025), and Qwen 3 235B IT Yang et al. (2025). We prompt each model using 8-shot examples from STEM for a given math standard and ask it to create a new MWP based on the same math standard (see Appendix J.4 for prompting details). Like other MWP generator works Ariyarathne et al. (2025); Christ et al. (2024); Sun et al. (2025), this prompting approach ensures fair comparisons across models by providing high-quality examples of MWPs in the expected format and keeping the prompt fixed across models.
@@
Table 2: Comparing LLMs for standards-aligned educational MWP generation. PPL is perplexity, BF1 is BERTScore F1, ASDIV BF1 compares each model’s MWPs to ASDIV’s 3rd-5th grade subset, length is average token length, readability is a composite of 8 metrics, and MaC is meets all criteria. Bold indicates the best performance in each metric (where applicable) and a * indicates the difference between the best open MaC performance and second best open MaC performance is significant at the p < 0.01 level. Standard errors, where applicable, are in parentheses.
@@ 4.1 Main Results
Using our Gemma 3 27B IT annotator, we evaluated each model’s MWPs for solvability, accuracy, educational appropriateness, and standards alignment. To reduce LLM calls, we annotate for all criteria together and have the LLM output if the MWP MaC along with its reasoning (see Appendix J.5 for an example). To validate our labels, our research team member with K-12 teaching experience and undergraduate education student labeled 100 randomly selected samples from our evaluation set. The annotators agreed with each other 76% of the time on MaC and had an average agreement rate of 75% with Gemma 3 27B IT for MaC. The Cohen’s $\kappa$ for MaC between the annotators (0.34) was nearly equivalent to the $\kappa$ for MaC between the annotators and Gemma 3 27B IT (0.30). Both Cohen’s Kappas suggest fair agreement, which is expected given MaC is a composite metric indicating agreement across four evaluation criteria and has unbalanced labels. Given the agreement rates and Cohen’s Kappas are nearly identical between experts and our automated annotator, we are confident in our labels.
Table 2 shows the MaC rate for each model. Our 30B model outperforms every baseline in MaC, while our 12B model outperforms Gemma 3 27B IT and nearly matches Qwen 3 30B IT, two larger and more capable models. The table also shows performance gaps between open and closed models and between smaller open and larger open models that our data/models eliminate or nearly eliminate, suggesting our data are sufficient for enabling open models to meet SOTA performance.
@@
Table 2 also shows PPL, BERTScore, and the average question and solution length and readability for each model’s MWPs. Our 12B model has the lowest PPL among all models and our 30B model has the lowest PPL among larger models, suggesting their MWPs are high quality. Our 12B model’s BERTScore is most similar to ASDIV’s within-dataset BERTScore among all models, while our 30B model ties with Qwen 3 235B IT in having a BERTScore closest to ASDIV’s among the open models excluding EDUMATH 12B. The BERTScores of our models’ MWPs compared to ASDIV are roughly equivalent to ASDIV’s within-dataset BERTScore, suggesting their MWPs are similar in diversity to human-written MWPs, while the MWPs from other models have higher BERTScores for this comparison. Our models’ average question lengths are most similar to those of STEM and ASDIV, suggesting their complexity, as measured by this proxy, is similar to teacher-written MWPs. Our models have the shortest solutions among open models, though the closed baselines have the shortest solutions among all models. Shorter solutions are preferable to longer, more verbose ones that may be daunting to read.
For readability, all models output questions with readability similar to STEM. For solution readability, all open models’ solutions have lower readability than STEM, while closed models exceed STEM’s readability, denoting a strength of open models. Overall, Table 2 suggests the EDUMATH models’ MWPs are high quality and similar to human-written MWPs. The alignment between our models’ performance across the automated evaluation criteria and higher rates of MaC further indicates our human evaluation criteria effectively identify high-quality MWPs.
@@
Although there are 38 math standard combinations, they can be grouped into combinations of one or more of 8 math topics: addition/subtraction, multiplication/division, fractions, decimals, perimeter/area/volume, patterns, elapsed time, and measurement conversion. Exploring MaC rates by topic can highlight the topics each model struggles with and succeeds at. Table 3 shows this comparison, while Appendix E shows further analysis of common error types for each model. EDUMATH 12B performs best at time problems and struggles the most with fraction problems, while EDUMATH 30B performs best at conversion problems and struggles the most with multiplication/division problems. EDUMATH 30B is the only model for which performance on each topic is above 90%, further highlighting its SOTA performance.
@@ 5 Evaluating MWPs with Students
While we show our models’ MWPs are high quality and can be customized to student interests (see Appendix F), it is critical to ensure students perform similarly on their MWPs relative to human-written MWPs and to identify if students prefer their customized MWPs to non-customized MWPs. To test this, 3rd-5th graders solved a worksheet with one MWP generated by EDUMATH 30B and aligned with topics they are interested in and one written by humans, either from existing data (K5Learning or ASDIV’s 3rd-5th grade subset) or written by our research team member with K-12 teaching experience if a worksheet’s math topic was not captured in existing data (see Appendix G.3 for an example). We also asked students which MWP they preferred and why. We tested these worksheets in two experimental settings at two elementary schools (see Appendix G.2 for school demographics). All worksheets were screened by teachers to ensure the MWPs were appropriate. At the first school, 4th graders (n = 82) solved one worksheet for each of four weeks where both the LLM-generated and human-written MWP were the same for all students. Each MWP was aligned with a standard students were currently learning and the LLM-generated MWP was based on a topic teachers identified as interesting to students. At the second school, 3rd-5th graders (n = 12) receiving services through a math interventionist completed one or two worksheets with the same human-written MWP for every student in each grade level and a customized LLM-generated MWP for each student in that grade level, depending on how many standards they were learning (see Appendix G.2 for details). To inform the customized MWPs, the interventionist surveyed students to identify topics they were interested in, resulting in deep customization for each student. These two experimental conditions allow us to test differential effects of customization at the class and individual level, while also testing our MWPs across all grade levels we study.
In this experiment (see Appendix G.1 for full results), students performed similarly on human-written and LLM-generated MWPs, further suggesting EDUMATH 30B’s MWPs are human quality. Students preferred the LLM-generated MWPs over the human-written MWPs in both conditions. In School 2 (customization at the individual level), every student but one preferred the LLM-generated MWPs. Students most often reported liking the LLM-generated MWP because they liked the topic it was about. These results offer compelling evidence that LLMs can support math education by providing customized MWPs that excite students.
@@ 6 Conclusion
We explore standards-aligned educational MWP generation and create the first dataset to train models for this task and the only one verified by teachers to have readable solutions. We also evaluate existing models at this task and fully release our annotations to aid future research. We demonstrate the quality of STEM and our annotated data by using them to develop the EDUMATH models, the first standards-aligned educational MWP generators. Our evaluations show EDUMATH 12B matches the performance of larger and more capable open models, while EDUMATH 30B sets a new SOTA. We also conduct the first study of custom LLM-generated MWPs with grade school students, finding they perform similarly on our MWPs relative to human-written MWPS, while preferring our MWPs. Our findings suggest LLMs can contribute to K-12 education by generating customized practice problems to reduce teacher burdens. Future research should develop models and data for other grade levels and types of math problems.
@@
Appendix A Model Development Pipeline
Figure 2 visually displays the process of developing the STEM dataset and EDUMATH models described in Section 3.
Figure 2: Visual of STEM dataset and EDUMATH models’ development process. Blue rounded rectangles denote LLMs, red cylinders denote data sources, purple rectangles denote model interventions/training processes, green funnels denote a dataset filtering process, directional arrows represent inputs/outputs, and MaC denotes meets all criteria.
Appendix B Annotator Directions and Demographics
Figures 3, 4, 5, and 6 display the directions presented to annotators for annotating MWPs for solvability, accuracy, educational appropriateness, and standards alignment, respectively. If an annotator answered "No" to solvability or any other criteria, they were instructed to select "NA" for the remaining criteria given that a failure to meet one of the criteria automatically means a MWP does not MaC. This saved the annotators time, as they did not need to unnecessarily annotate the remaining criteria. Table 4 displays annotator demographics and average time spent per MWP. Our human annotation study was approved by the institutional review board at our host university.